Find the angle between c and 2A if c=A+2B

1
Expert's answer
2021-10-24T18:27:23-0400
"\cos \alpha=\frac{(2\vec{A}\cdot(\vec{A}+2\vec{B}))}{2A(A+2B)}\\
\cos \alpha=\frac{A^2+2\vec{A}\cdot\vec{B}))}{A(A+2B)}"